A certain Force applied to mass m1 gives it an accelaration of 6m/s^2.the same force applied to mass m1 gives it an accelaration of 3m/s^2.if the two masses are fixed together and the same force is applied to the combination,accelaration of it would be

Grade:8

1 Answers

Arun
5 years ago
Dear student
 
1) m1 = F/6
2) m2 = F/3
Combining, m2/m1 = 2

a(2+1) = 6/(2+1) = 2 m/s²
 
Regards
Arun (askIITians forum expert)